Circular No. 8543

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 L4 (CHRISTENSEN) Eric J. Christensen, Lunar and Planetary Laboratory, reports his dis-

covery of a comet in the course of the Mt. Lemmon Survey on images taken with the 1.5-m reflector (discovery observation given below); the object displays a nearly stellar head with a faint tail ≈ 15′′ long in p.a. 250^{o} (as seen on four co-added 90-s images). Following posting on the ‘NEO Confirmation Page’, J. E. McGaha (Tucson, AZ, 0.62-m f/5.1 reflector) reports that his CCD images taken on June 14.3 UT show a compact coma of diameter ∼ 3′′ with a faint split tail 11′′ long in p.a. 260^{o} and 290^{o}.

2005 UT  α_{2000}  δ_{2000}  Mag.
June 13.35110  20^{h}11^{m}32.^{s}96  −19^{o}11′20′′.5  19.1

The available astrometry (including Spacewatch prediscovery observations back to June 3), the following elliptical orbital elements, and an ephemeris appear on MPEC 2005-L65.

T = 2005 Aug. 24.808 TT  ω = 24.^{o}789
e = 0.42406  Ω = 284.070 }2000.0
q = 2.36550 AU  i = 17.034
a = 4.10720 AU n^{o} = 0.118409 P = 8.32 years

COMET C/2005 K2 (LINEAR) Following news of this comet’s apparent outburst in brightness (cf.

IAUC 8540), several observers have detected what appears to be a splitting of the comet’s nucleus. M. Kidger, Instituto de Astrofisica de Canarias, writes that CCD images obtained by J. A. Reyes and S. Pastor (Murcia, Spain) show a secondary tailward condensation on June 12.88 (when it was 29′′ east and 25′′ north of the primary condensation, and 1.4 mag fainter) and 13.89 UT (29′′ east, 26′′ north; main condensation appears elongated). M. Meyer, Kelkheim, Germany, reports that images taken by M. Jäger and G. Rhemann (near Vienna, Austria) on June 12 show an apparent secondary condensation in the direction of the short tail; they add that this condensation shows more clearly on CCD images taken by D. Strange (Dorset, U.K., 0.50-m f/4 telescope) on June 13.98. G. Sostero and E. Guido report that remote CCD imaging using a 0.25-m f/3.4 reflector (near Mayhill, NM) reveals the following offsets from the primary and approximate magnitudes for the secondary: June 13.156, Δα = +1^{s}.91, Δδ = +22′′.2, R ∼ 17.1; June 14.152, Δα = +2^{s}.74, Δδ = +29′′.8, R ∼ 16.9.

2005 June 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green